# PAOLI: Pose-free Articulated Object Learning from Sparse-view Images

### v1

@@ Abstract
We present a novel self-supervised framework for learning articulated object representations from sparse-view, unposed images. Unlike prior methods that require dense multi-view observations and ground-truth camera poses, our approach operates with as few as four views per articulation and no camera supervision. To address the inherent challenges, we first reconstruct each articulation independently using recent advances in sparse-view 3D reconstruction, then learn a deformation field that establishes dense correspondences across poses. A progressive disentanglement strategy further separates static from moving parts, enabling robust separation of camera and object motion. Finally, we jointly optimize geometry, appearance, and kinematics with a self-supervised loss that enforces cross-view and cross-pose consistency. Experiments on the standard benchmark and real-world examples demonstrate that our method produces accurate and detailed articulated object representations under significantly weaker input assumptions than existing approaches.
@@ 1 Introduction
Interacting with articulated objects, objects that can change shape or position through joints, is a natural part of everyday human activity. Think of opening a cupboard door, sliding out a drawer, or lifting a lid. These objects are characterized by their ability to move in specific ways, often defined by a set of joints that allow for rotation or translation. Understanding how these objects operate is crucial for tasks such as manipulation, and control in robotics.
@@
Earlier research [35, 13, 39] tackled this challenge by using detailed ground-truth information such as 3D object shapes, articulation parameters, or part segmentations. While effective, this approach depends on high-quality 3D scans and labor-intensive manual annotations, which are both costly and difficult to scale. To address this, recent methods have shifted toward learning directly from raw, unlabeled data. Instead of relying on manual labels, these approaches capture the motion of articulated objects and use it to infer their structure and kinematics [22, 7, 10]. For instance, PARIS [22] and ArticulateYourNeRF [7] follow an analysis-by-synthesis approach, learn object representations from pairs of multi-view images, using Neural Radiance Fields (NeRF) [30] to model 3D geometry and appearance while separating moving parts from static ones. ArticulatedGS [10] improves over [22, 7] by modeling articulated objects through Gaussian Splatting (GS) [15], offering improved efficiency and accuracy. Despite these advances, existing methods still rely on dense observations, requiring around 100 paired views of the same object across two different articulated poses. Such requirements are rarely feasible in practical applications. Moreover, these approaches typically assume known camera poses, an assumption does not hold in many real-world scenarios.
@@
Figure 1: Given few unposed views of an articulated object over different articulated poses, PAOLI (our method) reconstructs its 3D geometry and appearance, while also estimating its part segmentation and joint axis of its moving parts (shown as arrows) in 3D. The reconstructed object can be rendered from novel viewpoints and articulated poses.
In this work, we propose a novel approach for learning articulated object representations from sparse-view, unposed images. This setting introduces two major challenges. First, with only a handful of views (e.g., four per articulation) and no camera information, reconstructing the object’s 3D structure becomes highly under-constrained. Second, recovering articulated motion and camera motion from images are tightly coupled, making it difficult to disentangle object motion from viewpoint changes.
To tackle these challenges, we introduce a novel framework, PAOLI, that learns to reconstruct articulated objects from sparse-view images without requiring camera poses. For the first challenge, we build on recent advances in sparse-view 3D reconstruction without camera supervision for static scenes. Specifically, we reconstruct the 3D shape of the object independently for each articulation [53], and then learn a deformation field that establishes dense correspondences across the reconstructed shapes. For the second challenge, we introduce a progressive disentanglement strategy: the procedure first identifies static components of the object before progressively isolating the moving parts. This step-by-step process enables a robust separation of camera motion from articulated motion. Finally, we jointly optimize geometry, appearance, and kinematics using a self-supervised loss that enforces consistency across different views and poses.
We evaluate our method on several benchmark datasets and show that it achieves accurate and detailed articulated object representations from sparse, unposed inputs, and significantly outperforming prior approaches that require dense multi-view data and ground-truth camera poses.
@@ 2 Related Work
Neural rendering [30, 15, 12, 45, 33] enables photorealistic novel view synthesis by rendering from learned 3D representations. More recently, 3D GS [15] has emerged as particularly compelling, offering efficient, high-quality representations [8, 5, 29, 18, 49]. Standard GS requires dense multi-view images with accurate poses. To relax this, some works use depth priors [21, 54, 6] or diffusion guidance [24, 16, 28], while others directly predict Gaussian parameters with feed-forward networks [52, 53, 40, 41, 2, 4]. Splatter Image [40] and LGM [41] handle single images, while MVSPlat [4] and PixelSplat [2] target sparse multi-view with known poses. FreeSplatter [53] and NoPoSplat [52] remove pose supervision by jointly predicting poses and Gaussians, but remain limited to static scenes without articulated motion.
@@
Reconstructing articulated objects is challenging due to coupled geometry and kinematics. Some methods focus on structure or motion alone [11, 17, 3, 9, 25, 14], while others use generative part-level graphs [23, 19]. More recent works jointly model geometry and appearance through optimization [32, 42, 47, 22, 7, 48, 27, 10]. PARIS [22] reconstructs parts and motions from dense multi-view scans with known poses; ArticulateYourNeRF [7] improves robustness via iterative segmentation and motion learning; while ArticulatedGS [10], ArtGS [27], and DTA [48] enhance quality with GS and depth priors. However, all rely on dense multi-view inputs and calibrated poses. Our approach instead reconstructs articulated objects from sparse, unposed views by leveraging recent feed-forward MVS advances.
@@ 3 Method
Assume that we are given two sets of images, each capturing an articulated object from sparse and unknown viewpoints. Within each set, the object remains in a fixed pose, while its articulated pose differs across the sets. Our objective is to reconstruct a digital twin of the object that includes both high-fidelity 3D geometry and appearance, and articulation parameters, enabling physically plausible dynamic rendering.
Our method consists of three main stages. First, we reconstruct the 3D geometry and estimate camera parameters for each set of sparse-view images independently using Gaussian representations. Next, we learn a deformation field to align the geometry between two sets. Based on the estimated deformation, we establish the initial dense correspondences and part segmentation. Finally, we jointly optimize the geometry, part segmentation, and articulation parameters to improve the overall reconstruction quality. An overview of our method is shown in Fig.˜2. Since the optimization of the geometry, part segmentation, and articulation parameters is intertwined and requires good initialization, the first two stages are crucial for the success of the final joint optimization.
@@ 3.1 Initialization of 3D Gaussian Splats
Since the standard 3D GS [15] typically requires many input views for accurate reconstruction, we use FreeSplatter [53], a feed-forward neural network trained to reconstruct 3D Gaussian splats from sparse views. As FreeSplatter is designed for static scenes, we apply it independently to each image set to obtain initial Gaussian splats and camera parameters.
Despite its ability to work in the sparse setting, the Gaussians estimated by FreeSplatter are not accurately pixel-aligned with the input images, resulting in suboptimal camera extrinsic estimates. Additionally, the estimated camera intrinsics often differ between the two sets. Hence, we use their average as the final intrinsic parameters. To improve the pixel alignment, we refine the extrinsic parameters $\Pi^{s}$ and $\Pi^{t}$ by directly optimizing them on the reconstructed Gaussian splats over their respective multi-view set, minimizing photometric loss via rendering. This two-step camera refinement yields better alignment between the estimated camera parameters and the 3D geometry, providing a better initialization for the subsequent steps.
3.2 Gaussian Alignment via Deformation Fields
A central challenge in building digital twins of articulated objects from pose-free images lies in aligning the canonical pose across different observation sets. We frame this alignment as estimating the deformation between the Gaussian representations $G^{s}$ and $G^{t}$. A straightforward strategy is to optimize the mean and rotation of each Gaussian point independently. However, this naive approach breaks local rigidity, since it allows unconstrained updates of individual parameters, whereas true object motion should remain rigid within each part.
To address this, we introduce a feed-forward deformation network, $\mathcal{F}_{\text{deform}}$, which predicts rigid transformations for Gaussian points in $G^{s}$ using their positional encodings, producing a deformed target representation $\hat{G}^{t}$. By constraining the network to generate low-frequency deformations, we enforce local rigidity while still capturing global motion. Specifically, $\mathcal{F}_{\text{deform}}$ maps the positional encoding of Gaussian means to a rigid transformation consisting of a rotation, translation, and scaling factor:
@@ 3.2 Gaussian Alignment via Deformation Fields
where $\hat{G}^{t}=\{\hat{\mathbf{g}}_{i}^{t}\}_{i=1}^{N^{s}}$ represents the deformed source Gaussians, with each point $\hat{\mathbf{g}}^{t}$ obtained by applying the transformation predicted by $\pi=\mathcal{F}_{\text{deform}}(\bm{\mu})$ to the corresponding source point $\mathbf{g}^{s}\in G^{s}$, where $\pi$ is a similarity transformation containing rotation $\mathbf{R}$, translation $\mathbf{t}$, and scaling $s$. Specifically, to obtain the deformed target Gaussian points $\hat{\mathbf{g}}^{t}=(\hat{\bm{\mu}}^{t},\hat{\mathbf{q}}^{t},\mathbf{s}^{s},o^{s},\mathbf{c}^{s})$, we first compute the transformation for each point $\mathbf{g}^{s}=(\bm{\mu}^{s},\mathbf{q}^{s},\mathbf{s}^{s},o^{s},\mathbf{c}^{s})$ using Eq.˜1, and apply the transformation to the means and rotation of the source Gaussian points. With $\hat{\bm{\mu}}^{t}=s\mathbf{R}\bm{\mu}^{s}+\mathbf{t}$ and $\hat{\mathbf{q}}^{t}=\mathbf{q}^{s}\otimes\mathbf{q}(\mathbf{R})$, we can obtain the deformed target Gaussian points $\hat{\mathbf{g}}^{t}=(\hat{\bm{\mu}}^{t},\hat{\mathbf{q}}^{t},\mathbf{s}^{s},o^{s},\mathbf{c}^{s})$, where $\otimes$ denotes quaternion multiplication and $\mathbf{q}(\mathbf{R})$ converts the rotation matrix $\mathbf{R}$ to its quaternion representation. For simplicity, we denote this process as $\hat{\mathbf{g}}=\pi\circ\mathbf{g}$.
@@ 3.3 Joint Optimization
To obtain an initial segmentation of object parts and their corresponding rigid transformations, we leverage TEASER [50], a robust pose estimation algorithm for point correspondences. We apply TEASER to the set of correspondences between the deformed source and the target Gaussians. Since TEASER was originally designed for static scenes, it identifies the largest rigidly moving component, typically the part with greatest volume, as the inlier set, while treating the remaining points as outliers.
We then organize these sets into a tree structure: the inlier points form the root node, and the outliers are assigned to leaf nodes. TEASER is applied recursively to the leaf nodes to estimate part-level similarity transformations. This iterative process progressively adds parts to the tree, with the root representing the dominant component and leaves corresponding to smaller articulated parts. For simplicity, we describe the two-part case, though the method naturally extends to multiple parts. The result is a set of similarity transformations, $\pi_{r}$ and $\pi_{l}$, for the root and leaf parts, along with segmentation labels $p^{s}\in\{r,l\}$ for each Gaussian point in the source state.
Next we proceed to the joint optimization of the Gaussian splats, part segmentation, and articulation parameters. We initialize the learnable Gaussian splats $\bar{G}^{s}=\{\bar{\mathbf{g}}^{s}_{i}\}_{i=1}^{N^{s}}$ with $\bar{\mathbf{g}}^{s}=(\bm{\mu}^{s},\mathbf{q}^{s},\mathbf{s}^{s},o^{s},\mathbf{c}^{s},p^{s})$ and treat $\pi_{r}$ and $\pi_{l}$ as learnable parameters to refine the rigid transformation. To obtain the deformed Gaussian splats, we apply the transformations to each Gaussian point based on its segmentation label: $\pi_{r}\circ\mathbf{g}$ if $p^{s}=r$, or $\pi_{l}\circ\mathbf{g}$ if $p^{s}=l$. Since the initial estimation is not perfect, we also apply a displacement term $\delta\bm{\mu}$ to the means of the deformed Gaussian points to compensate for non-rigidity. To simplify the notation, we use an auxiliary function $f(\mathbf{g},\pi,\delta\bm{\mu})$ that first applies the transformation $\pi$ to $\mathbf{g}$ and then applies the displacement $\delta\bm{\mu}$ to the mean of the transformed Gaussian. This yields the deformed Gaussian splats $\bar{G}^{t}=\{\bar{\mathbf{g}}^{t}|\bar{\mathbf{g}}^{t}=f(\bar{\mathbf{g}}^{s},\pi_{p^{s}},\delta\bm{\mu}),\bar{\mathbf{g}}^{s}\in\bar{G}^{s}\}$, with the $\delta\bm{\mu}$ denoting the displacement for the means.
Since the initial segmentation and and transformation for each part is imperfect. We often end up with detached parts after the deformation. To deal with this, we refine the estimated $\pi_{r}$ and $\pi_{l}$ with the initial segmentation labels in the first phase of optimization using both the Chamfer distance between the transformed $\bar{G}^{t}$ and $G^{t}$, and the rendering loss computed in both the source and target states. We define the joint loss as:
@@
and our optimization objective is to minimize the joint loss:
@@
Note that in this phase, $\bm{\delta}\mu$ is set to zero.
In the second phase of the optimization, we optimize the Gaussian splats to improve the geometric accuracy, especially to handle occlusion between source and target states. We also allow the displacement term $\delta\bm{\mu}$ to be optimized to compensate for non-rigidity caused by imperfect transformation estimation or geometry mismatch, but still penalize the displacement of the Gaussian splats with $\mathcal{L}_{\text{rigid}}=\sum_{i=1}^{N}\left\|\delta\bm{\mu}_{i}\right\|_{1}$. For better optimization stability suggested in ArticulateYourNeRF [7], we update the segmentation labels heuristically. Newly split or duplicated points inherit the segmentation labels from their parent points. To correct points with incorrect initial segmentation, we check those points with large $\delta\bm{\mu}$, which indicates high non-rigidity, and update their segmentation labels following this rule:
@@
with $\bar{\bm{\mu}}^{p}$ for the means of $\bar{\mathbf{g}}^{p}=\pi_{p}\circ\bar{\mathbf{g}}^{s},p\in\{r,l\}$. And in this phase, our optimization objective is to minimize the joint loss with the regularization term:
@@
and the update of the segmentation labels is:
$\min_{\pi_{r},\pi_{l}}\mathcal{L}_{\text{joint}}$ (8)
In the final phase, we focus on improving visual quality and articulation estimation. We update the $\pi_{r}$, $\pi_{l}$, and $\mathbf{c}$ of the Gaussian splats without the displacement term $\delta\bm{\mu}$ for pure rigid transformation supervised by photo loss with the original input images as:
@@
In practice, we use 500 iterations for the first phase, 2500 iterations for the second phase, and 2000 iterations for the final phase. After 5000 iterations at total, we decompose the articulation parameters into axis $\mathbf{a}\in\mathbb{R}^{3}$, angle $\theta\in\mathbb{R}$, and pivot point $\mathbf{p}\in\mathbb{R}^{3}$ for revolute joints, while only axis $\mathbf{a}$ and magnitude $d\in\mathbb{R}$ for prismatic joints.
3.4 Extension to Multiple Moving Parts

### v2

@@ Abstract
We present a methodology to model articulated objects using a sparse set of images with unknown poses.
Current methods require dense multi-view observations and ground-truth camera poses. Our approach operates with as few as four views per articulation and no camera supervision. Our central insight is to first solve a robust correspondence and alignment problem between unaligned reconstructions, before part motions can be analyzed. We first reconstruct each articulation independently using recent advances in sparse-view 3D reconstruction, then learn a deformation field that establishes dense correspondences across poses. A progressive disentanglement strategy further separates static from moving parts, enabling robust separation of camera and object motion. Finally, we optimize geometry, appearance, and kinematics jointly with a self-supervised loss that enforces cross-view and cross-pose consistency. Experiments on the standard benchmark and real-world examples demonstrate that our method produces accurate and detailed articulated object representations under significantly weaker input assumptions than existing approaches.
@@ 1 Introduction
Automating human capabilities, such as reasoning about and interacting with pivoting doors, sliding drawers, opening boxes, remains an open problem. Practical solutions for modeling such articulated objects—rigid objects coupled via joints that allow rotation or translation— from video is beneficial to downstream applications such as robotic manipulation. Current approaches require dense, multi-view images across articulation poses, making them impractical. In this work, we show that articulated object modeling is possible with only four unposed images per articulation state, vastly reducing capture requirements.
@@
Early approaches [42, 18, 48] rely on rich supervision such as 3D object shapes, articulation parameters, or part segmentations. While effective, this approach depends on high-quality 3D scans and labor-intensive manual annotations, which are both costly and difficult to scale. Recent methods reduce supervision by learning from unlabeled multiple views; rather than relying on manual labels, these approaches infer object structure and kinematics from observed motion [30, 10, 16]. For instance, PARIS [30] and ArticulateYourNeRF [10] follow an analysis-by-synthesis approach, learn object representations from pairs of multi-view images, using Neural Radiance Fields (NeRF) [38]. These models jointly recover 3D geometry and appearance while separating moving parts from static ones. ArticulatedGS [16] further improves efficiency and accuracy by replacing NeRF with Gaussian Splatting (GS) [21].
Despite these advances, the state of the art still relies on dense, paired observations, requiring around 100 paired views of the same object across two articulation states. Such dense capture is rarely practical. Moreover, these methods assume that all capture lie in a shared, pre-calibrated coordinate system, i.e., that the object’s static base is perfectly aligned across articulation states via known camera poses. In practice, this level of calibration is costly and rarely achievable, severely limiting deployment.
@@
Figure 1: Given few unposed views of an articulated object across different articulated poses, PAOLI (our method) reconstructs its 3D geometry and appearance, while estimating part segmentation and joint axis of its moving parts (shown as arrows) in 3D. The reconstructed object can be rendered from novel viewpoints and articulation configurations.
We instead study articulated object reconstruction and modeling from unposed and sparse views, significantly more realistic yet fundamentally harder setting. Traditional 3D reconstruction methods such as structure-from-motion (SFM) [45, 46] or neural-based methods [53, 55] assume static scenes and cannot directly handle object motion. Treating each articulation state independently bypasses this issue but produces reconstructions in different coordinate frames, creating a severe misalignment problem. Articulated reasoning becomes impossible without first establishing a reliable 3D-to-3D correspondence between these unaligned shapes.
A common approach is to match sparse keypoints. However, keypoints are unreliable under large articulations, on textureless surfaces, and provide insufficient coverage for dense part segmentation. This makes them unsuitable for the experimental setting we target.
We introduce PAOLI, a methodology that learns dense correspondences via a optimized deformation field. Leveraging sparse view 3D reconstruction [62, 14], we first obtain independent GS reconstructions for each articulation state. A dense deformation field is optimized to warp the source model into alignment with the target state by enforcing 3D geometric consistency and 2D photometric appearance. Dense correspondences between the original source model and the target model are then established by using 1-Nearest-Neighbor (1-NN) to match the deformed source model to the target model. Finally to separate rigid parts, we introduce an iterative robust estimation algorithm that clusters inliers from the correspondence set using a pose estimator [59], thereby isolating static and moving components and recovering their kinematic parameters.
In summary, our main contributions are:
  1. 1.
    Problem formulation: We are the first to address articulated object reconstruction from sparse, unposed images, identifying the 3D-to-3D correspondence problem between unaligned reconstructions as the central challenge.
  2. 2.
    Dense correspondence via deformation: We propose a robust correspondence detector, using a dense deformation field, which is optimized with 3D and 2D losses. This approach outperforms keypoint-based correspondence methods and other adapted baselines.
  3. 3.
    A new customized solution: We demonstrate the first complete pipeline to successfully recover full articulated models including geometry, parts, and kinematics from sparse, and unposed data.
These contributions demonstrate that a robust solution to this correspondence problem is achieved not by sparse keypoints or pre-trained models, but by a dense deformation field optimized at test-time for each object, leveraging a combination of 3D geometry and 2D rendering losses.
@@ 2 Related Work
Neural rendering [38, 21, 17, 54, 41] enables photorealistic novel view synthesis by rendering from learned 3D representations. More recently, 3D GS [21] has emerged as particularly compelling, offering efficient, high-quality representations [12, 8, 37, 25, 58]. Standard GS requires dense multi-view images with accurate poses. To relax this, some works use depth priors [29, 63, 9] or diffusion guidance [32, 22, 36], while others directly predict Gaussian parameters with feed-forward networks [61, 62, 49, 50, 4, 6]. Splatter Image [49] and LGM [50] handle single images, while MVSPlat [6] and PixelSplat [4] target sparse multi-view with known poses. FreeSplatter [62] and NoPoSplat [61] remove pose supervision by jointly predicting poses and Gaussians, but remain limited to static scenes without articulated motion. UFV-splatter [14] improves robustness to image viewpoints by finetuning FreeSplatter with images from unfavorable views, which enhances reconstruction and camera pose estimation.
@@
Reconstructing articulated objects is challenging due to coupled geometry and kinematics. Some methods focus on structure or motion alone [19, 23, 5, 13, 33, 20], while others use generative part-level graphs [31, 26]. More recent works jointly model geometry and appearance through optimization [40, 51, 56, 30, 10, 57, 35, 16]. DITTO [19] as one of the prioneer work that aims to model two-part articulated objects with point cloud input data and assumes aligned coordinate frame with pre-trained feed-forwrd model. PARIS [30] reconstructs parts and motions from dense multi-view scans with known poses; ArticulateYourNeRF [10] improves robustness via iterative segmentation and motion learning; while ArticulatedGS [16], ArtGS [35], and DTA [57] enhance quality with GS and depth priors. However, all rely on dense multi-view inputs and calibrated poses. Our approach instead reconstructs articulated objects from sparse, unposed views by leveraging recent feed-forward MVS advances.
3D-3D Correspondence and Matching
Finding 3D-to-3D correspondences is a fundamental problem. Approaches based on sparse keypoints, from classical descriptors like FPFH [44] to modern learned detectors [28, 64]. Recent registration methods for 3D Gaussian Splatting (GS) like GaussReg [3], proposed to establish coarse 3D keypoint matching between GS, which still struggles with accurate registration in 3D data, making them unsuitable for dense articulated motion analysis. A more robust alternative is to learn dense correspondences via deformation fields [15, 11, 2, 24, 7]. These approaches, however, typically rely on feed-forward, pre-trained models to predict deformation and are designed for meshes or point clouds, not 3D GS. In contrast, our method introduces a test-time, optimization-based framework to learn a dense deformation field specifically for GS, allowing us to solve the articulated motion problem. We then use a robust iterative algorithm to segment these dense correspondences into distinct rigid parts and their kinematics.
@@ 3 Method
We address the task of reconstructing articulated objects, defined as a composition of rigid parts connected by 1-DoF revolute or prismatic joints. Given two sets of images, each capturing an articulated object from sparse and unknown viewpoints (with a fixed pose within each set but different poses across sets), our objective is to reconstruct a ”digital twin” of the object. This model includes high-fidelity 3D geometry, appearance, and articulation parameters, enabling physically plausible dynamic rendering.
Our method consists of three main stages. First, we reconstruct the 3D geometry and estimate camera parameters for each image set independently. Next, we learn a deformation field to align the geometry and establish initial dense correspondences. Finally, we jointly optimize the geometry, part segmentation, and articulation parameters. An overview is shown in Fig. 2. Since the final optimization is intertwined and requires good initialization, the first two stages are crucial.
@@ 3.1 Initialization of 3D Gaussian Splats
Since FreeSplatter is designed for static scenes, we apply it independently to the source set $\mathcal{I}^{s}$ and the target set $\mathcal{I}^{t}$. This process yields our two initial 3D Gaussian Splatting models, $G^{s}$ and $G^{t}$, each in its own arbitrary coordinate system. This step also provides initial camera parameters (both extrinsics and intrinsics) for each set. We then compute a single, averaged intrinsic matrix from the two sets and use this fixed intrinsic parameter for all subsequent optimization steps.
However, the initial Gaussians and extrinsic estimates are not perfectly pixel-aligned. To refine this, we freeze the GS and optimize only the extrinsic parameters ($\Pi^{s},\Pi^{t}$) for each set by minimizing a photometric rendering loss. This refinement provides a better initialization for the subsequent steps.
3.2 Gaussian Correspondences via Deformation Fields
A central challenge in building digital twins of articulated objects from pose-free images lies in aligning the canonical pose across different observation sets. As argued in our related work, a common approach for recovering pose is to match sparse keypoints [28, 64]. However, this is often unreliable for dense analysis and struggles with texture-less surfaces. A more robust alternative is to find a dense alignment. A naive strategy would be to optimize the mean and rotation of each Gaussian point independently, but this breaks local rigidity, since it allows unconstrained updates of individual parameters, whereas true object motion should remain rigid within each part.
To address this, we learn a per-scene deformation field, parameterized by a lightweight feed-forward network, $\mathcal{F}_{\text{deform}}$. This network is optimized from scratch for each object and predicts rigid transformations for Gaussian points in $G^{s}$ using their positional encodings, producing a deformed target representation $\hat{G}^{t}$. By constraining the network to generate low-frequency deformations, we enforce local rigidity while still capturing global motion. Specifically, $\mathcal{F}_{\text{deform}}$ maps the positional encoding of Gaussian means to a rigid transformation consisting of a rotation, translation, and scaling factor:
@@ 3.2 Gaussian Correspondences via Deformation Fields
where $\hat{G}^{t}=\{\hat{\mathbf{g}}_{i}^{t}\}_{i=1}^{N^{s}}$ represents the deformed source Gaussians, with each point $\hat{\mathbf{g}}^{t}$ obtained by applying the transformation predicted by $\pi=\mathcal{F}_{\text{deform}}(\bm{\mu})$ to the corresponding source point $\mathbf{g}^{s}\in G^{s}$, where $\pi$ is a similarity transformation containing rotation $\mathbf{R}$, translation $\mathbf{t}$, and scaling $s$. Specifically, to obtain the deformed target Gaussian points $\hat{\mathbf{g}}^{t}=(\hat{\bm{\mu}}^{t},\hat{\mathbf{q}}^{t},\mathbf{s}^{s},o^{s},\mathbf{c}^{s})$, we first compute the transformation for each point $\mathbf{g}^{s}=(\bm{\mu}^{s},\mathbf{q}^{s},\mathbf{s}^{s},o^{s},\mathbf{c}^{s})$ using Eq. 1, and apply the transformation to the means and rotation of the source Gaussian points. With $\hat{\bm{\mu}}^{t}=s\mathbf{R}\bm{\mu}^{s}+\mathbf{t}$ and $\hat{\mathbf{q}}^{t}=\mathbf{q}^{s}\otimes\mathbf{q}(\mathbf{R})$, we can obtain the deformed target Gaussian points $\hat{\mathbf{g}}^{t}=(\hat{\bm{\mu}}^{t},\hat{\mathbf{q}}^{t},\mathbf{s}^{s},o^{s},\mathbf{c}^{s})$, where $\otimes$ denotes quaternion multiplication and $\mathbf{q}(\mathbf{R})$ converts the rotation matrix $\mathbf{R}$ to its quaternion representation. For simplicity, we denote this process as $\hat{\mathbf{g}}=\pi\circ\mathbf{g}$.
Pipeline Corr. Foldchair Fridge Laptop Oven Scissor Stapler USB Washer Blade Storage
Motion Ang Err $\downarrow$ AGS[16] - 62.083 36.568 86.058 78.129 62.083 54.613 88.677 87.454 59.562 75.744
Ours GaussReg[3] failed failed failed failed failed failed failed failed failed failed
Ours FM-Dense[2] 0.974 13.267 failed failed 5.209 failed 12.603 35.926 3.716 49.799
Ours Ours 0.095 0.914 1.020 2.365 0.839 1.337 1.872 1.737 0.878 0.525
Pos Err $\downarrow$ AGS[16] - 0.032 0.124 1.805 0.252 0.032 0.149 1.209 0.231 - -
Ours GaussReg[3] failed failed failed failed failed failed failed failed - -
Ours FM-Dense[2] 0.729 0.658 failed failed 0.394 failed 0.281 0.417 - -
Ours Ours 0.180 0.007 0.106 0.042 0.058 0.196 0.099 0.157 - -
Motion Dist $\downarrow$ AGS[16] - 70.900 75.548 46.925 29.381 70.900 64.865 61.955 33.345 1.275 0.213
Ours GaussReg[3] failed failed failed failed failed failed failed failed failed failed
Ours FM-Dense[2] 13.590 25.261 failed failed 3.652 failed 11.285 17.509 0.089 0.192
Ours Ours 0.861 2.279 0.788 2.654 1.516 0.930 1.877 2.926 0.017 0.118
Rendering PSNR $\uparrow$ AGS[16] - 9.349 11.294 13.787 13.871 13.481 12.898 15.856 17.644 24.543 14.333
Ours GaussReg[3] failed failed failed failed failed failed failed failed failed failed
Ours FM-Dense[2] 17.897 19.256 21.161 19.356 23.276 19.002 22.517 20.632 25.741 20.900
Ours Ours 24.659 26.920 26.491 23.434 28.823 26.532 27.016 27.416 29.041 25.493
SSIM $\uparrow$ AGS[16] - 0.652 0.722 0.809 0.723 0.846 0.831 0.842 0.877 0.962 0.765
Ours GaussReg[3] failed failed failed failed failed failed failed failed failed failed
Ours FM-Dense[2] 0.889 0.916 0.923 0.893 0.953 0.930 0.821 0.933 0.970 0.896
Ours Ours 0.940 0.963 0.955 0.931 0.967 0.953 0.956 0.966 0.974 0.927
Geometry CD $\downarrow$ AGS[16] - 1188.824 717.780 893.149 244.862 1188.824 1599.261 467.633 268.377 90.288 231.355
Ours GaussReg[3] failed failed failed failed failed failed failed failed failed failed
Ours FM-Dense[2] 4.814 6.399 failed failed 0.501 failed 8.935 25.515 2.789 15.599
Ours Ours 0.856 2.051 0.610 6.449 0.264 1.216 1.430 1.615 1.788 6.443
@@ 3.3 Joint Optimization
To obtain an initial segmentation and transformations, we leverage TEASER++ [59], a robust pose estimation algorithm. For clarity, we first describe the common two-part (root/leaf) scenario (see Sec. 3.4 for multi-part). We apply TEASER++ to the correspondences from Sec 3.2. It finds the largest rigid component (inliers) as the root $r$, treating the remaining outliers as the leaf $l$. This yields initial transformations $\pi_{r},\pi_{l}$ and part labels $p^{s}$.
Both the initial geometry from FreeSplatter and the part-pose estimates from TEASER++ are coarse approximations. Therefore, a joint optimization is essential to mutually refine the geometry, segmentation, and articulation parameters. To do this without the optimization getting stuck in a bad local minimum, we introduce a three-phase staged refinement strategy. This coarse-to-fine approach first stabilizes the large-scale rigid motions (Phase 1), then refines the underlying geometry and segmentation (Phase 2), and finally polishes the visual appearance (Phase 3).
We define our full learnable model and its parameters. We initialize the learnable Gaussian splats $\bar{G}^{s}=\{\bar{\mathbf{g}}^{s}_{i}\}_{i=1}^{N^{s}}$, where each $\bar{\mathbf{g}}^{s}=(\bm{\mu}^{s},\mathbf{q}^{s},\mathbf{s}^{s},o^{s},\mathbf{c}^{s},p^{s})$ is augmented with its initial part label $p^{s}$. We also treat the rigid transformations $\pi_{r}$ and $\pi_{l}$ as learnable parameters. To compensate for non-rigidity and initial estimation errors, we introduce a learnable, per-Gaussian displacement term $\delta\bm{\mu}$. To simplify the notation, we define an auxiliary function $f(\bar{\mathbf{g}}^{s}_{i},\delta\bm{\mu}_{i},\pi)$. This function first selects the appropriate rigid transformation $\pi_{p^{s}}$ from the set $\pi$ (using the part label $p^{s}$ contained within $\bar{\mathbf{g}}^{s}_{i}$), applies it, and then adds the displacement $\delta\bm{\mu}_{i}$ to the transformed mean. This yields the final deformed Gaussian splats $\bar{G}^{t}=\{\bar{\mathbf{g}}^{t}_{i}|\bar{\mathbf{g}}^{t}_{i}=f(\bar{\mathbf{g}}^{s}_{i},\delta\bm{\mu}_{i},\pi),\bar{\mathbf{g}}^{s}_{i}\in\bar{G}^{s}\}$.
Phase 1: Rigid Pose Refinement The goal of this phase is to stabilize the coarse poses from TEASER++. We apply two constraints: the source Gaussian geometry $\bar{G}^{s}$ is frozen, and the non-rigid displacement $\delta\bm{\mu}$ is set to zero. Using the fixed initial segmentation $p^{s}$, we only optimize the rigid transformations $\pi=\{\pi_{r},\pi_{l}\}$ by minimizing a joint loss:
@@
The optimization objective for this phase is:
@@
Phase 2: Geometry and Occlusion Refinement With the rigid poses stabilized in Phase 1, we now address the inaccuracies in the underlying geometry and the initial part segmentation. In this phase, we unfreeze the parameters in $\bar{G}^{s}$ and activate the learnable displacement term $\delta\bm{\mu}$. This per-Gaussian term aims at compensating for any non-rigid errors, which typically caused by incorrect part segmentations. We encourage the displacement to be sparse by penalizing it with $\mathcal{L}_{\text{rigid}}=\sum_{i=1}^{N}\left\|\delta\bm{\mu}_{i}\right\|_{1}$.
Crucially, this displacement term also serves as a heuristic to correct mis-segmented points. As suggested in [10], we periodically update segmentation labels $p^{s}$. We assume points with a large $\delta\bm{\mu}$ are mis-assigned, and we re-assign their label $p$ based on the smallest rigid alignment error:
@@
with $\bar{\bm{\mu}}^{p}$ for the means of $\bar{\mathbf{g}}^{p}=\pi_{p}\circ\bar{\mathbf{g}}^{s},p\in\{r,l\}$.
The full optimization objective minimizes the joint loss plus the rigidity penalty, updating all parameters (geometry, transformations, and displacements) simultaneously:
@@
Phase 3: Pose and Appearance Refinement The previous phases refined the geometry and produced an as-rigid-as-possible part segmentation. In this final phase, we lock this geometry (freezing $\bar{G}^{s}$ and setting $\delta\bm{\mu}=\mathbf{0}$) to focus exclusively on finetuning the per-part pose estimation ($\pi_{r},\pi_{l}$) and visual quality (colors $\mathbf{c}$). This is supervised purely by the photometric loss against the original input images:
@@
In practice, we use 500 iterations for the first phase, 2500 iterations for the second phase, and 2000 iterations for the final phase. Finally, we decompose the articulation parameters into axis $\mathbf{a}\in\mathbb{R}^{3}$, angle $\theta\in\mathbb{R}$, and pivot point $\mathbf{p}\in\mathbb{R}^{3}$ for revolute joints, while only axis $\mathbf{a}$ and magnitude $d\in\mathbb{R}$ for prismatic joints.
3.4 Modeling Multiple Moving Parts